OpenStudy (anonymous):

Solve s^2 + 10s + 21 = 0 Solution is s=

OpenStudy (saifoo.khan):

s = -3 s = -7

OpenStudy (anonymous):

s^2 + 10s + 21 = 0 s^2+ 7s+3s+21=0 s(s+7) + 2(s+7) = 0 (s+2)(s+7) = 0 either s+2 = 0 or s+7=0 hence s=-2 or s=-7